OpenStudy (cicilybailey14):

Fan+Medal...solve the pattern! Then find the next two numbers in the pattern. 4, 6, 9, 13.5, ... 20, 27 15.5, 19 20.25, 30.375 20, 27.5

OpenStudy (cicilybailey14):

@AnswerMyQuestions @ganeshie8 @iGreen. @pooja195

OpenStudy (anonymous):

4->6=4+(4/2) 6->9=6+(6/2) 9->13.5=9+(9/2) 13.5->X=13.5+(13.5/2) X->Y=X+X/2

OpenStudy (anonymous):

I think it divides the number by 2, then it adds 3...

OpenStudy (anonymous):

*Then it multiples 3.

OpenStudy (anonymous):

4 / 2 = 2. 2 x 3 = 6. 6 / 2 = 3. 3 x 3 = 9. 9 / 2 = 4.5. 4.5 x 3 = 13.5.

OpenStudy (anonymous):

Divide by 2, then multiply to 3.

OpenStudy (anonymous):

@Cicilybailey14 Divide 13.5 / 2.

OpenStudy (cicilybailey14):

6.75

OpenStudy (anonymous):

Multiply that to 3.

OpenStudy (anonymous):

It's funny because both patterns are the same :3

OpenStudy (cicilybailey14):

okay...20.25

OpenStudy (anonymous):

Yes, so what is your answer?

OpenStudy (anonymous):

Since only 1 choice has 20.25 :p

OpenStudy (cicilybailey14):

Lol yeah, :F Thx!

OpenStudy (anonymous):

Np.